The electric potential at any point is given
by ¢= xy(z²— 3xy). Calculate the electric field E at that point.
1
Expert's answer
2019-05-27T10:52:41-0400
"\mathbf{E} = -\nabla \phi = - (\partial_x \phi, \partial_y \phi, \partial_z \phi)"


"\partial_x \phi = \partial_x (xyz^2 - 3x^2 y^2) = yz^2 -6xy^2""\partial_y\phi =xz^2 - 6x^2 y""\partial_z \phi = 2xyz"


"\mathbf{E} = (6xy^2 - yz^2, 6x^2 y - xz^2, -2xyz)"